I need 500 mL of 0.2 M acetate, pH 5.0. I have solid anhydrous sodium acetate (MW=82 g/mol^-1) and a solution of 1M acetic acid. Describe how to make the buffer.
Answer #1

total moles of buffer = 0.5 L x 0.2 mole / L = 0.1

x (acetic acid ) + y (sodium acetate) = 0.1 -----------------> (1)

pH = pKa + log (y / x)

5.0 = 4.76 + log (y/x )

log (y/x ) = 0.24

y / x = 1.74

y = 1.74 x -----------------------> 2

by solving 1 and 2

x = 0.0365

y = 0.0635

moles of acetic acid = 0.0365

molarity of aceitic acid given = 1 M

molarity = moles / volume

1   = 0.0365 / volume

volume = 0.0365 / 1 = 0.0365 L

volume = 36.5 mL

moles of sodium acetate = 0.0635

mass of sodium acetate = 0.0635 x 82 = 5.21 g

to prepare this buffer : 5.21 g soild sodium acetate and 36.5 mL acetic acid must be taken